# Semi-supervised Underwater Image Enhancement Using A Physics-Aware Triple-Stream Network

### v5

@@ III-B PATS-UIENet
Figure 2: The architecture of the proposed PATS-UIENet, which comprises three individual streams, namely, D-Stream, B-Stream and A-Stream, to estimate the degradation parameters of the revised IFM [2].
@@
Red Channel Tuner (RCT). Inspired by the channel attention mechanism [52], we propose a compact RCT module in order to adaptively emphasize the information contained in the red channel. This module first uses a convolutional layer to extract basic features from the input image, and then use the Global Average Pooling (GAP) to obtain a global representation. A fully-connected layer and the Sigmoid activation function are further used to transform this representation to a tuning weight. The weight is finally used to scale the red channel of the image.
@@
D-Stream. D-stream is designed to estimate the direct signal transmission map of an underwater image. According to Eq. (3), the transmission map varies at the locations of different pixels. In this case, the network should be able to generate fine-grained representations. Hence, we build the D-Stream using a CNN-based encoder-decoder network, due to its strong ability to learn local characteristics, which helps preserve edges and fine details. The encoder comprises five consecutive blocks, denoted as $Enc\_1$ to $Enc\_5$. As a result, multi-scale feature maps are derived using the encoder. Given $Enc\_5$ is used as the bottleneck, the decoder contains four blocks, denoted as $Dec\_1$ to $Dec\_4$, symmetrically. Skip connection is used to pass the feature maps produced by an encoder block to the corresponding decoder block at the same level, which is useful for restoring the fine-grained spatial structure. The use of multi-scale features and skip connections enhances the ability of the network to restore the spatial structure of the scene. The output of the last decoder block is fed into a convolutional layer. Three transmission maps are produced in terms of different color channels. In essence, D-Stream is specifically focused on modeling the directly transmitted light, which normally contains the spatial structure of the scene.
B-Stream. B-stream is responsible for modeling the degradation caused by the backscattering in underwater environments. Although it shares the same CNN-based architecture as D-Stream for training stability, B-stream differs in functionality and processing strategy. Unlike direct signals, backscatter signals exhibit strong global interference and require a context-aware estimation. To this end, we pass the output of the first encoder block $Enc\_1$ through a Patchify processing and feed it into the A-Stream, leveraging the global modeling capability of Transformer to enhance the estimation accuracy. Moreover, B-Stream is connected with the A-Stream via Residual Communication Modules (RCMs), enabling bidirectional feature exchange. Compared to the D-Stream, B-Stream is focused on modeling the global degradation rather than recovering the local scene content.
A-Stream. Ambient light reflects the overall luminance in the underwater scene and is generally spatially homogeneous. To effectively model this global characteristic, we employ a Transformer-based design instead of CNNs, because Transformer [53] is well-suited for capturing long-range dependencies and global characteristics through the self-attention mechanism. A-Stream consists of five Transformer blocks, denoted as $Trans\_1$ to $Trans\_5$, and includes a dedicated “Ambient” token to specifically learn the ambient light parameter. Additionally, A-Stream is coupled with the B-Stream through the RCM, allowing it to incorporate local features and mitigate the lack of spatial detail inherent in Transformer. A-Stream is distinguished from the other streams by its global feature learning process and Transformer-based architecture, which complements the local characteristic modeling of the CNN-based streams.
@@
Residual Communication Module (RCM). According to the revised IFM [2], the degradation can be considered as the outcome caused by both the transmission and the ambient light. In this situation, the backscatter signal transmission estimation steam and the ambient light estimation stream will learn some common features. Therefore, the feature exchange between both the streams is likely to boost the training of each stream. Recently, the complementary action of CNNs and Transformers has been shown [54]. We are inspired to propose the RCM for the sake of bridging the two streams. Specifically, the tokens produced by a block in the A-stream are folded into a set of feature maps while the feature maps generated by a block of the encoder of the B-stream are downsampled to the shape of these maps. Then both sets of maps are concatenated along the channel dimension and are fed into a convolutional layer. Finally, the resultant maps are split into two sets along the channel dimension. Each set is added with the original feature maps of the related stream.
III-C IFM-Inspired Semi-supervised Learning Framework
@@
To overcome the challenge of insufficient real-world underwater images, we adopt a semi-supervised learning framework (see Fig. 3), inspired by the revised IFM [2] theory, embedding physical constraints directly into the learning process to effectively leverage unlabeled data. This framework includes a bi-directional supervised scheme and an unsupervised scheme. The bi-directional scheme contains a forward-enhancement module, which learns from reference images and uses the estimated parameters of the revised IFM to perform UIE, and a backward-degradation module, which uses the estimated parameters to degrade the reference images to their real-world degraded counterparts. As a result, the bi-directional supervised scheme can learn a more accurate estimation of the parameters using limited real-world data than that learned using a supervised method with a large number of synthetic data. Also, the generalization of the model trained is stronger than that trained using a supervised method with limited real-world data. On the other hand, the unsupervised scheme exploits the unlabeled real-world underwater images that we collect. A second set of images with different degrees of degradation are generated from these images based on the revised IFM [2]. Two sets of parameters can be estimated using both sets of images, respectively. Each set is utilized as the reference for the other set because their IFMs are related.
@@
Figure 3: The proposed semi-supervised learning framework, which comprises a bi-directional supervised learning scheme and an unsupervised learning scheme, used for training our PATS-UIENet.
@@
$\mathcal{L}_{fwd}=||(J^{c}(x)-\hat{J}^{c}(x))||_{2}^{2}.$ (4)
Due to the ill-posed nature of the revised IFM, however, only $\mathcal{L}_{fwd}$ may be insufficient for ensuring the validity of the estimated parameters. Therefore, we propose a backward-degradation learning module which degrades clear reference images using the estimated parameters to their degraded counterparts $\hat{I}^{c}(x)$. The supervision over the backward-degradation is achieved by minimizing the loss function:
$\mathcal{L}_{bwd}=||(I^{c}(x)-\hat{I}^{c}(x))||_{2}^{2}.$ (5)
@@
Unsupervised Learning. To leverage unlabeled underwater images, we adopt an unsupervised learning scheme based on the physical modeling capability of the revised IFM [2] rather than using generative networks [34, 19], which are difficult to train and often produce unstable results. This scheme adaptively constructs pairs of images with different levels of degradation from the unlabeled data.
Given an unlabeled real-world degraded image $I^{c}_{1}(x)$, the revised IFM can be expressed as:
$I^{c}_{1}(x)=J^{c}_{1}(x)t_{D1}^{c}+(1-t_{B1}^{c}(x))A^{c}_{1}.$ (8)
A new degraded image $I^{c}_{2}(x)$ can be purposely derived according to:
$I^{c}_{2}(x)=\alpha I^{c}_{1}(x)+(1-\alpha)(1-t_{B1}^{c}(x))A^{c}_{1},$ (9)
@@
$I^{c}_{2}(x)=J^{c}_{1}(x)(\alpha t_{D1}^{c}(x))+(1-t_{B1}^{c}(x))A^{c}_{1}.$ (10)
Eq. (10) also satisfies the revised IFM but presents a more severe degradation than the $I^{c}_{1}(x)$, generated from the same underlying clear image $J^{c}_{1}(x)$, because $\alpha t_{D1}^{c}(x)<t_{D1}^{c}(x)$ always holds true. Therefore, the parameters $\hat{t}_{D2}^{c}(x)$ and $\hat{t}_{B2}^{c}(x)$ estimated using the PATS-UIENet from image $I^{c}_{2}(x)$ will be close to $\alpha\hat{t}_{D1}^{c}(x)$ and $\hat{t}_{B1}^{c}(x)$, respectively, where $\hat{t}_{D1}^{c}(x)$ and $\hat{t}_{B1}^{c}(x)$ are the parameters estimated from $I^{c}_{1}(x)$. Correspondingly, two loss functions are defined as:
$\mathcal{L}_{D}=||\hat{t}_{D2}^{c}(x)-\alpha\hat{t}_{D1}^{c}(x)||_{2}^{2},$ (11)
$\mathcal{L}_{B}=||\hat{t}_{B1}^{c}(x)-\hat{t}_{B2}^{c}(x)||_{2}^{2}.$ (12)
Similar to Eq. (6), the loss function for the ambient light $\hat{A}^{c}$ estimated by the A-stream is defined as:
$\mathcal{L}_{A\text{-}unsup}=||(I^{c}(x)*G-\hat{A}^{c}||_{2}^{2}.$ (13)
@@
$\mathcal{L}_{unsup}=\mathcal{L}_{D}+\mathcal{L}_{B}+\mathcal{L}_{A\text{-}unsup}+\lambda_{3}\mathcal{L}_{gw},$ (15)
@@ IV-A Baselines
We compared the proposed method with two prior-based approaches which were developed on top of the IFM [1], including UDCP [4] and DHL [10], three prior-based methods which did not consider the IFM, including Histogram Equalization (HE) [9], HLRP [15] and MMLE [14]. We also compared our method with seven supervised learning approaches, including WaterNet [22], UColor [23], FUnIE-GAN [20], PUIE-Net (MP) [25], U-Transformer [21], URanker [26] and CCMSRNet [27], an unsupervised learning method, i.e., USUIR [24], and three semi-supervised learning methods, including DDFormer [29], Semi-UIR [28] and UWFormer [30].
IV-B Data Sets
@@
We conducted a series of experiments using five data sets, including four publicly available real-world underwater data sets, i.e., UIEB [22], SUIM-E [31], RUIE [56] and SQUID [10], and a synthetic data set, namely, EUVP [20]. The UIEB data set originally contains 890 pairs of degraded images, in which each reference image was selected from the results produced by applying 12 algorithms to a degraded image. Ten pairs of UIEB [22] images were discarded because the degraded images in these pairs had been included in the rest pairs. In total, we only used 880 pairs of UIEB images.
The SUIM-E [31] data set comprises 1,525 pairs of degraded underwater images and a testing set of 110 labeled images. Three subsets are comprised of the RUIE [56] data set, including UIQS, UCCS and UHTS, which include 3,630, 300 and 300 unlabeled images, respectively. The SQUID [10] data set comprises 114 underwater images captured using a stereo camera, divided into the Michmoret, Katzaa, Nachsholim and Satil subsets. We utilized 2,185 images from the Underwater Scenes subset of the EUVP [20] data set.
Two different PATS-UIENet models were trained for the real-world and synthetic testing images, referred to as $PATS\text{-}UIENet_{Real}$ and $PATS\text{-}UIENet_{Syn}$, respectively. Regarding both the models, we randomly selected 720 pairs from the 880 pairs of UIEB [22] images and 2,000 pairs from the Underwater Scenes subset of the EUVP [20] data set, respectively, which were used as the training images in the supervised learning scheme. We collected 1,934 unlabeled real-world underwater images, which cover diverse underwater scenes, and used these images in the unsupervised learning scheme for both the models.
@@
Among the rest of the UIEB [22] data set, 80 pairs of images were randomly selected as the validation set, while the remaining pairs were used as the testing set which is referred to as Test-U80. The UIEB data set also consists of 60 challenging images without references, which were used as the second testing set, referred to as Test-C60. A third testing set was obtained from the testing set of the SUIM-E data set [31], which contains 110 pairs of labeled images and is named as Test-S110. For the RUIE [56] data set, the UCCS was utilized as the fourth testing set, denoted as Test-UCCS. Fifty-three SQUID [10] images captured using the camera at the right hand side were randomly selected and were used as the fifth testing set, namely, Test-R53. The remaining 185 images in the Underwater Scenes subset of the EUVP [20] data set were used as the sixth testing set, referred to as Test-Scenes.
IV-C Evaluation Metrics
@@
For the labeled testing data, we used Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) [57] and Learned Perceptual Image Patch Similarity (LPIPS) [58] to assess the quality of an enhanced image with regard to the reference image. Regarding the unlabeled testing data, we used the Underwater Image Fidelity (UIF) [59] and Multi-scale Image Quality (MUSIQ) [60] metrics. When the SQUID [10] data set was tested, the Average Reproduction Angular Error ($\bar{\psi}$) [10] was used to evaluate the quality of color restoration. Since transmission maps are related to the scene distance, we used the Pearson Correlation Coefficient (PCC) calculated between the estimated transmission maps and the ground-truth depth map to assess the performance of transmission estimation [10].
IV-D Implementation Details
@@
We implemented the proposed method using Pytorch and conducted experiments on Ubuntu 20.04 with a GeForce RTX 3090 graphics processing unit. During the training process, the images were resized to a resolution of $256\times 256$ pixels. The number of filters in the RCT was set to 16. For the $Enc\_1$ to $Dec\_1$ in the D-Stream and B-Stream, the numbers of filters were set to 64, 128, 256, 512, 512, 256, 128, 64 and 64 in turn. Regarding the A-Stream, the dimension of each token was 384 and the number of attention heads in the MHSA was 6. We trained the PATS-UIENet using the AdamW [61] optimizer. The learning rate and batch size were set to 1e-4 and 12, respectively. Four weighting factors, including $\lambda_{1},\lambda_{2},\lambda_{3}$ and $\lambda_{unsup}$, were set to 0.1, 0.005, 1 and 0.1, respectively. We first trained the PATS-UIENet using the bi-directional supervised scheme for 50 epochs as a warm-up stage. Then the semi-supervised scheme was used to train it for 1000 epochs. When the semi-supervised learning process was carried out, in particular, the unsupervised scheme was performed for 30 iterations after the bi-directional supervised scheme was conducted for each epoch.
V Experimental Results

### v6

@@ 3 Methodology
Figure 2: The architecture of the proposed PA-UIENet, which comprises two individual streams: A-Stream and T-Stream, to estimate the degradation parameters of the IFM.
3.1 PA-UIENet
As illustrated in Fig.2, the PA-UIENet contains an encoder-decoder style T-stream and a Transformer-based A-stream, used to estimate transmission maps and the ambient light respectively. Since the red channel usually encounters the more severe attenuation than the blue and green channels [1, 38], the utilization of this channel is key to color restoration. Therefore, a simple Red Channel Tuner (RCT) module is designed to adaptively emphasize the red channel before an image is fed into the T-stream. The output of the first encoder block Enc_1Enc\_1 in the T-stream is processed by Patchify and the result is sent to the A-stream. A Residual Communication Module (RCM) connects the two blocks of the encoder of the T-stream and the A-stream at the same level, to fulfill the bi-directional feature exchange. Both the transmission maps and the ambient light, estimated using the T-stream and A-stream respectively, are fed into the IFM. As a result, the enhanced underwater image is produced.
@@ 3.1 PA-UIENet
Red Chanel Tuner (RCT). Inspired by the channel attention mechanism [39], we propose a compact RCT module in order to adaptively emphasize the information contained in the red channel. This module first uses a convolutional layer to extract basic features from the input image, and then use the Global Average Pooling (GAP) to obtain a global representation. A fully-connected layer and the Sigmoid activation function are further used to transform this representation to a tuning weight. The weight is finally used to scale the red channel of the image.
@@
T-Stream. According to Eq. (1), the transmission map varies at different pixel locations. In this case, the network should be able to generate fine-grained representations. Hence, we build the T-stream using a CNN-based encoder-decoder network. The encoder comprises five blocks, denoted as Enc_1Enc\_1 to Enc_5Enc\_5, and the Max Pooling operation is applied to each block. As a result, multi-scale feature maps are derived using the encoder. Given Enc_5Enc\_5 is used as the bottleneck, the decoder contains four blocks, denoted as Dec_1Dec\_1 to Dec_4Dec\_4, symmetrically. Bilinear interpolation is used to upsample the feature maps within each block. Both the encoder and decoder blocks share the same structure, which consists of two ConvConv-BNBN-ReLUReLU operations. Skip connection is used to pass the feature maps produced by an encoder block to the corresponding decoder block at the same level, which is useful for restoring the fine-grained spatial structure. The output of the last decoder block is fed into a convolutional layer. A set of transmission maps are produced in terms of different color channels.
A-Stream. The ambient light describes the luminance of the scene. Since this parameter does not encode the information of a specific object, it can be treated as homogeneous. We are motivated to design the A-stream using a set of Transformer [40] blocks for the sake of leveraging the self-attention mechanism to extract the global information. Five blocks, denoted as Trans_1Trans\_1 to Trans_5Trans\_5, are comprised of this stream. Each block contains a self-attention layer and a two-layer Multi-layer Perceptron (MLP) unit, in which two residual connections are used respectively. Since the input of the Transformer is a sequence of visual tokens , the feature maps of Enc_1Enc\_1 are processed using Patchify which includes the Flatten and Linear Projection operations. Considering the estimation of the ambient light does not require local characteristics and the quadratic complexity of self-attention, we downsample these maps by a factor of 16 to reduce the computational demand. The result is a set of tokens. A special token, namely, the Ambient token, is also adopted, to specifically learn the ambient light parameter. All the tokens are fed into Trans_1Trans\_1. The Ambient token produced by the last block Trans_5Trans\_5 is sent to a linear layer. Finally, the estimation of the ambient light is achieved.
@@
Residual Communication Module (RCM). According to the IFM, the degradation can be considered as the outcome caused by both the transmission and the ambient light. In this situation, the T-stream and A-stream will learn some common features. Therefore, the feature exchange between both the streams is likely to boost the training of each stream. Recently, the complementary action of CNNs and Transformers has been shown [41]. We are inspired to propose the RCM for the sake of bridging the two streams. Specifically, the tokens produced by a block in the A-stream are folded into a set of feature maps while the feature maps generated by a block of the encoder of the T-stream are downsampled to the shape of these maps. Then both sets of maps are concatenated along the channel dimension and are fed into a convolutional layer. Finally, the resultant maps are split into two sets along the channel dimension. Each set is added with the original feature maps of the related stream.
3.2 IFM-Inspired Semi-Supervised Learning Framework
@@
To overcome the challenge of insufficient real-world underwater images, we adopt a semi-supervised learning framework (see Fig.3), inspired by the IFM theory. This framework includes a bi-directional supervised scheme and an unsupervised scheme. The bi-directional scheme contains a forward-enhancement module, which learns from reference images and uses the estimated IFM parameters to perform UIE, and a backward-degradation module, which uses estimated parameters to degrade the reference images to their real-world degraded counterparts. As a result, the bi-directional supervised scheme can learn a more accurate estimation of the parameters using limited real-world data than that learned using a supervised method with a large number of synthetic data. Also, the generalization of the model trained is stronger than that trained using a supervised method with limited real-world data. On the other hand, the unsupervised scheme exploits the unlabeled real-world underwater images that we collect. A second set of images with different degrees of degradation are generated from these images based on the IFM. Two sets of parameters can be estimated using both sets of images respectively. Each set is utilized as the reference for the other set because their IFMs are related.
@@
Figure 3: The proposed semi-supervised learning framework, which comprises a bi-directional supervised learning scheme and an unsupervised learning scheme, used for training our PA-UIENet.
@@
[…]
Due to the ill-posed nature of the IFM, however, only $\mathcal{L}_{fwd}$ may be insufficient for ensuring the validity of the estimated parameters. Therefore, we propose a backward-degradation learning module which degrades clear reference images using the estimated parameters to their degraded counterparts $\hat{I}^{c}(x)$. The supervision over the backward-degradation is achieved by minimizing the loss function:
[…]
@@
Unsupervised Learning. To leverage unlabeled underwater images, generative networks [21, 15] were used. However, they were difficult to train and produced unstable results. Since unsupervised learning methods [23, 24] are able to directly learn from degraded images without requiring reference images, we are motivated to adopt an unsupervised learning scheme on top of the IFM.
Given an unlabeled real-world degraded image $I^{c}_{1}(x)$, the IFM can be expressed as:
$I^{c}_{1}(x)=J^{c}_{1}(x)t^{c}_{1}(x)+(1-t^{c}_{1}(x))A^{c}_{1}.$ (6)
As performed in [23], a new degraded image $I^{c}_{2}(x)$ can be purposely derived according to:
$I^{c}_{2}(x)=\alpha I^{c}_{1}(x)+(1-\alpha)A^{c}_{1},$ (7)
@@
$I^{c}_{2}(x)=J^{c}_{1}(x)(\alpha t^{c}_{1}(x))+(1-\alpha t^{c}_{1}(x))A^{c}_{1}.$ (8)
Eq. (8) also satisfies the underwater IFM but presents the more severe degradation than the $I^{c}_{1}(x)$, generated from the same underlying clear image $J^{c}_{1}(x)$, because $\alpha t^{c}_{1}(x)<t^{c}_{1}(x)$ always holds true. Therefore, the parameters $\hat{t}^{c}_{2}(x)$ and $\hat{A}^{c}_{2}$ estimated using the PA-UIENet from image $I^{c}_{2}(x)$ will be close to $\alpha\hat{t}^{c}_{1}(x)$ and $\hat{A}^{c}_{1}$ respectively, where $\hat{t}^{c}_{1}(x)$ and $\hat{A}^{c}_{1}$ are the parameters estimated from $I^{c}_{1}(x)$ [23]. Correspondingly, two loss functions are defined as:
[…]
@@
$\mathcal{L}_{unsup}=\mathcal{L}_{T}+\mathcal{L}_{A\text{-}unsup}+\lambda_{3}\mathcal{L}_{gw},$ (12)
@@ 4 Experimental Settings
Baselines. We compared the proposed method with four prior-based approaches which were developed on top of the IFM, including DCP [2], UDCP [3], GDCP [4] and DHL [8], two prior-based methods which did not consider the IFM, including Histogram Equalization (HE) and Retinex [11], and two supervised deep learning approaches, including WaterNet [16] and UColor [17].
@@
Data Sets. We conducted a series of experiments on five testing sets derived from four publicly available real-world underwater data sets: UIEB [16], SUIM-E [18], RUIE [43] and SQUID [8]. The UIEB data set originally contains 890 pairs of degraded images, in which each reference image was selected from the results produced by applying 12 algorithms to a degraded image. Ten pairs of UIEB [16] images were discarded because the degraded images in these pairs had been included in the rest pairs. Only 880 pairs of UIEB images were used in this study. The SUIM-E data set comprises 1,525 pairs of degraded underwater images and a testing set of 110 labeled images. Three subsets are comprised of the RUIE data set, including UIQS, UCCS and UHTS, which include 3,630, 300 and 300 unlabeled images, respectively. The SQUID data set comprises 114 underwater images captured using a stereo camera, divided into the Michmoret, Katzaa, Nachsholim and Satil subsets.
@@
Regarding the supervised learning, we randomly selected 720 pairs from the 880 pairs UIEB [16] images and used these as the training set. For the unsupervised learning, we collected 1,934 unlabeled real-world underwater images, which cover diverse underwater scenes. In total, our training set contained 2,654 images. Among the rest of the UIEB [16] data set, 80 pairs of images were randomly selected as the validation set, while the remaining pairs were used as the testing set which is referred to as Test-U80. The UIEB data set also consists of 60 challenging images without reference, which were used as the second testing set, referred to as Test-C60. A third testing set was obtained from the testing set of the SUIM-E data set [18], which contains 110 pairs of labeled underwater images and is named as Test-S110. For the RUIE [43] data set, the UCCS was utilized as the fourth testing set, denoted as Test-UCCS. Besides, 53 SQUID [8] images captured using the camera at the right hand side were randomly selected and were used as the fifth testing set, namely, Test-R53.
@@
Evaluation Metrics. For the labeled testing data, we used both Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) [44] to assess the quality of an enhanced image with regard to the reference image. Regarding the unlabeled testing data, we used the Underwater Image Quality Measure (UIQM) [45] and Underwater Color Image Quality Evaluation (UCIQE) [46] metrics. When the SQUID [8] data set was tested, the Average Reproduction Angular Error ($\bar{\psi}$) [8] was used to evaluate the quality of color restoration. Since transmission maps are related to the scene distance, we used the Pearson Correlation Coefficient (PCC) calculated between the estimated transmission maps and the ground-truth depth map to assess the performance of transmission estimation [8].
@@
Implementation Details. We implemented our method using Pytorch and conducted experiments on Ubuntu 20.04 with a GeForce RTX 3080 graphics processing unit. During the training process, the images were resized to the resolution of $256\times 256$ pixels. The number of filters in the RCT was 16. For the $Enc_{1}$ to $Dec_{1}$ in the T-stream, the numbers of filters were 64, 128, 256, 512, 512, 256, 128, 64 and 64 in turn. Regarding the A-stream, the dimension of each token was 384 and the number of attention heads in the MHSA was 6. We trained the PA-UIENet using the AdamW [47] optimizer. The learning rate and batch size were set to 1e-4 and 6 respectively. Four weighting factors, including $\lambda_{1},\lambda_{2},\lambda_{3}$ and $\lambda_{unsup}$ were set to 0.001, 0.005, 10 and 0.001 respectively. We first trained the PA-UIENet using the bi-directional supervised scheme for 3,000 iterations as a warm-up stage. Then the semi-supervised scheme was used to train it for 150,000 iterations. When the semi-supervised learning was carried out, in particular, the unsupervised scheme was performed for 30 iterations after the bi-directional supervised scheme was conducted for every 120 iterations.
5 Experimental Results